Mondi Company produces party boxes that are sold in bundles of 1000 boxes. The market is highly competitive, with boxes currently selling for R100 per thousand. The company has a total and marginal cost curve given by:

TC = 3,000,000 + 0.001Q2

MC = 0.002Q

Q is measured in thousand box bundles per year. [5]

a. Determine Mondi's profit maximizing quantity.

b. Calculate if the firm is earning a profit or a loss?

c. Based on the analysis above, should Mondi Company operate or shut down in the shortrun?


1
Expert's answer
2021-10-22T08:18:56-0400

Solution:

a.). Profit maximizing quantity for a competitive firm: P = MR = MC

P = 100

MC = 0.002Q

Set: P = MC

100 = 0.002Q

Q = 50,000

Mondi’s profit maximizing quantity = 50,000 units

 

b.). Profit = TR – TC

TR = P x Q = 100 x 50,000 = 5,000,000

TC = 3,000,000 + 0.001Q2 = 3,000,000 + 0.001(50,0002) = 3,000,000 + 2,500,000 = 5,500,000

Profit = 5,000,000 – 5,500,000 = (500,000)

Loss = (500,000)

The firm is earning a loss of 500,000

 

c.). Derive average variable cost:

AVC ="\frac{0.001Q^{2} }{Q}" = 0.001Q = (0.001 "\times" 50,000) = 50

P>AVC

In the short run, if a firm is operating at a loss, it must decide to either continue with operations or temporarily shut down. The shutdown rule states that in the short run a firm should continue to operate if the price exceeds average variable costs.

Since the firm’s price is above the average variable costs, then it should not shut down but continue with operations.